### v2

@@ Abstract
In this work, we systematically study the interactions of the $S$-wave $D^{(*)}\bar{B}^{(*)}$ systems within the framework of chiral effective field theory in heavy hadron formalism. We calculate the $D^{(*)}\bar{B}^{(*)}$ effective potentials up to next-to-leading order, explore the bound state formations, and investigate the $D^{(*)}\bar{B}^{(*)}$ scattering properties such as scattering rate, scattering length, and effective range. Our results show that all $I=1$ $D^{(*)}\bar{B}^{(*)}$ potentials are repulsive, preventing the formation of bound states, while the $I=0$ potentials are generally attractive. Specifically, we get two important observations: first, the shallow bound state is more likely to exist in the $D\bar{B}[I(J^{P})=0(0^{+})]$ system than in the $D\bar{B}^{*}[I(J^{P})=0(1^{+})]$ system; second, $D^{*}\bar{B}^{*}[I(J^{P})=0(0^{+})]$ and $D^{*}\bar{B}^{*}[I(J^{P})=0(1^{+})]$ systems possess relatively large binding energies and positive scattering lengths, which suggests strong bound state formations in these channels. So the attractions in the $D^{*}\bar{B}^{*}[I=0]$ systems are deeper than those in the $D\bar{B}^{(*)}[I=0]$ systems, thus we strongly recommend the future experiment to search for the $D^{*}\bar{B}^{*}[I=0]$ tetraquark systems. We also investigate the coupled-channel effects on the $J=0,1$ systems and conclude that the inclusion of the coupled channels introduces small but visible influences. In addition, we also investigate the dependencies of the $D\bar{B}^{(*)}$ binding energies on the contact low-energy coupling constants (LECs).
@@ I Introduction
The study of exotic hadrons has become a vibrant frontier in particle physics, offering profound insights into the nonperturbative regime of strong interaction. In recent years, the LHCb Collaboration has made groundbreaking discoveries that have significantly advanced our understanding of multiquark states. In 2020, LHCb reported the observation of two resonances, $X_{0}(2900)$ and $X_{1}(2900)$, in the $D^{-}K^{+}$ invariant mass spectrum of the $B^{+}\to D^{+}D^{-}K^{+}$ decay process LHCb:2020bls ; LHCb:2020pxc . These states, with quantum numbers $J^{P}=0^{+}$ and $J^{P}=1^{+}$, respectively, were identified as the first tetraquark candidates with the exotic quark content $ud\bar{s}\bar{c}$. This discovery was followed in 2021 by the observation of the double-charm tetraquark state $T^{+}_{cc}$ in the $D^{0}D^{0}\pi^{+}$ invariant mass spectrum LHCb:2021vvq ; LHCb:2021auc . Shortly thereafter, LHCb reported two additional open heavy-flavor tetraquark candidates, $T_{c\bar{s}0}^{a}(2900)^{++}$ and $T_{c\bar{s}0}^{a}(2900)^{0}$, in the $D_{s}^{+}\pi^{\pm}$ final states of the $B^{+}\to D^{-}D^{+}_{s}\pi^{+}$ and $B^{0}\to\bar{D}^{0}D^{+}_{s}\pi^{+}$ decays LHCb:2022sfr ; LHCb:2022lzp . These states are believed to have the quark content $c\bar{s}u\bar{d}$.
@@
Doubly heavy dimeson systems have attracted significant theoretical interest due to their potential to form stable or quasi-stable molecular states. For instance, Li et al. explored the possibility of deuteron-like molecular states in $D^{(*)}D^{(*)}$, $\bar{B}^{(*)}\bar{B}^{(*)}$, and $D^{(*)}\bar{B}^{(*)}$ systems, identifying several promising candidates Li:2012ss . QCD sum rules have been applied to study exotic open-flavor tetraquark states such as $bc\bar{q}\bar{q}$, $bc\bar{s}\bar{s}$, $qc\bar{q}\bar{b}$, and $sc\bar{s}\bar{b}$, revealing that some of these states lie below the $D^{(*)}B^{(*)}$ and $D_{s}^{(*)}B_{s}^{(*)}$ thresholds, making them susceptible to decay via fall-apart mechanisms Chen:2013aba . The color-magnetic interaction model has also been used to calculate mass splittings for $qq\bar{Q}\bar{Q}$ tetraquark states Luo:2017eub , while the molecular picture has been employed to investigate the binding energies of $T_{cc}^{+}$ and its bottom and strange partners Deng:2021gnb . Additionally, the complex scaling method (CSM) has been utilized to study doubly heavy tetraquark bound and resonant states Wu:2024zbx .
The stability of doubly heavy tetraquarks, particularly those involving bottom and charm quarks, remains a topic of active debate. While some studies suggest that states such as $bb\bar{u}\bar{d}$, $bb\bar{u}\bar{s}$, and $bb\bar{d}\bar{s}$ are stable against strong decay Eichten:2017ffp , others argue that bottom-charm tetraquarks may not be stable Silvestre-Brac:1993zem ; Ebert:2007rn ; Park:2018wjk ; Lu:2020rog ; Braaten:2020nwp ; Kim:2022mpa ; Song:2023izj . The production mechanisms and detection prospects for these states at colliders have also been explored, with estimates suggesting significant production cross sections for double-bottom and mixed bottom-charm tetraquarks at the LHC Ali:2018xfq .
Recent lattice QCD studies have provided further insights into the properties of doubly heavy tetraquarks. Meinel et al. found no evidence for QCD-stable $\bar{b}\bar{c}ud$ tetraquark states Meinel:2022lzo , while Alexandrou et al. and Padmanath et al. reported shallow bound states for isoscalar $\bar{b}\bar{c}ud$ systems Alexandrou:2023cqg ; Padmanath:2023rdu . These findings, along with the identification of sub-threshold poles in the $D\bar{B}$ scattering amplitude Radhakrishnan:2024ihu , significantly advanced our understanding of the weak decay behaviors of $D\bar{B}$ molecules Liu:2024lmv .
@@ II.1 Lagrangians at the leading order
where $D_{a}$, $D_{b}$, $E_{a}$, and $E_{b}$ are independent LECs, $\tau^{\alpha}$ represents the Pauli matrix in isospin space. The heavy meson doublet $(D,D^{*})$ or $(\bar{B},\bar{B}^{*})$ is described by the $H$ field in the heavy hadron formalism Xu:2017tsr ; Xu:2021vsi ; Wang:2018atz :
@@
where $v=(1,0,0,0)$ is the four-velocity of the heavy mesons. For the $D^{*}\bar{B}^{*}$ system, the scattering amplitudes receive contributions from the one-pion exchange interactions at LO. The Lagrangian for the LO $D^{*}D^{*}\pi$ vertex or $\bar{B}^{*}\bar{B}^{*}\pi$ vertex is given by Xu:2017tsr ; Xu:2021vsi ; Wang:2018atz
@@
where $\Delta$ in the last term represents the mass splitting between $D$ and $D^{*}$ (or $\bar{B}$ and $\bar{B}^{*}$), $\Gamma_{\mu}$ is the chiral connection, $u_{\mu}$ is the axial vector current, and $\xi=\exp(i\phi/2f)$. $f$ is the bare pion decay constant, and the pion field $\phi$ is
@@ II.2 Lagrangians at next-to-leading order
$\mathcal{L}_{4H}^{(2,h)}=$ $D_{a}^{h}\mathrm{Tr}[H\gamma_{\mu}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\bar{H}]\mathrm{Tr}(\chi_{+})+D_{b}^{h}\mathrm{Tr}[H\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}(\chi_{+})$
$+E_{a}^{h}\mathrm{Tr}[H\gamma_{\mu}\tau^{a}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\tau_{a}\bar{H}]\mathrm{Tr}(\chi_{+})+E_{b}^{h}\mathrm{Tr}[H\gamma_{\mu}\gamma_{5}\tau^{a}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\gamma_{5}\tau_{a}\bar{H}]\mathrm{Tr}(\chi_{+}),$ (6)
$\mathcal{L}_{4H}^{(2,v)}=$$\{D_{a1}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}(v\cdot D\bar{H})]\mathrm{Tr}[H\gamma^{\mu}\bar{H}]+D_{a2}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}\bar{H}]\mathrm{Tr}[(v\cdot DH)\gamma^{\mu}\bar{H}]$
$+D_{a3}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}(v\cdot D\bar{H})]+D_{a4}^{v}\mathrm{Tr}[((v\cdot D)^{2}H)\gamma_{\mu}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\bar{H}]$
@@
$+E_{b1}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}\gamma_{5}\tau^{a}(v\cdot D\bar{H})]\mathrm{Tr}[H\gamma^{\mu}\gamma_{5}\tau_{a}\bar{H}]+\cdots\}+\mathrm{H.c.},$ (7)
$\mathcal{L}_{4H}^{(2,q)}=$$\{D_{1}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}(D^{\nu}\bar{H})]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}\bar{H}]+D_{2}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[(D^{\nu}H)\gamma_{\nu}\gamma_{5}\bar{H}]$
$+D_{3}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}(D^{\nu}\bar{H})]+D_{4}^{q}\mathrm{Tr}[(D^{\mu}D^{\nu}H)\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}\bar{H}]$
$+E_{1}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}\tau^{a}(D^{\nu}\bar{H})]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}\tau_{a}\bar{H}]+\cdots\}+\mathrm{H.c.},$ (8)
@@
The low-energy constants appearing in the above equations are split into finite and infinite parts. The infinite parts are used to cancel the divergences of the loop contributions at NLO. The finite parts also contribute but are not determined here due to the lack of data input. Therefore, we just discard them.
After the scattering amplitudes for the $D^{(*)}\bar{B}^{(*)}$ systems are calculated using the above Lagrangians, the effective potentials in momentum space can be obtained via the relation:
@@
where the factor $-1/4$ comes from the Breit approximation. To understand the interactions and further investigate whether the $D^{*}\bar{B}^{(*)}$ systems can form stable molecular states, we need the potentials in coordinate space. Fourier transformations of the momentum-space potentials allow for their derivation:
@@
Simultaneously, according to the effective potentials in momentum space, we can also solve the Lippmann-Schwinger equations (LSE) to calculate the partial-wave two-body scattering amplitudes $T$:
@@
Considering the $S$-wave, in the low-energy limit the $k\rm{cot}\delta_{0}(k)$ can be expanded in powers of $k^{2}$:
@@
where $a$ is the scattering length and $r_{0}$ is the effective range. The ellipsis represents higher-order terms in $k$. The $S$-wave scattering cross-section is proportional to the phase shift:
@@ III.1 $D\bar{B}$ systems
Next, we consider NLO contributions illustrated in Fig. 1. There are two types of NLO diagrams. The diagrams in Fig. 1 represent one-loop corrections to the LO contact interaction, and the corresponding amplitudes are given by:
@@
Using the Lagrangians Eq. (3), the amplitudes for the TPE diagrams in Fig. 1 can be written as
$\mathcal{M}_{c1.1}^{(2)}=$ $-4\frac{1}{f^{4}}[A_{1}(q_{0}^{2}J_{0}^{F}+J_{22}^{F})-A_{15}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})-A_{51}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})+A_{5}(q_{0}^{2}J_{0}^{F}+2q_{0}^{2}J_{11}^{F}$
$+q_{0}^{2}J_{21}^{F}+J_{22}^{F})],$ (25)
$\mathcal{M}_{c1.2}^{(2)}=$ $-4i\frac{g_{2}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}-(d-1)q_{0}J_{21}^{T}$
$+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]\},$ (26)
$\mathcal{M}_{c1.3}^{(2)}=$ $-4i\frac{g_{2}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}-(d-1)q_{0}J_{21}^{T}$
$+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]\},$ (27)
$\mathcal{M}_{c1.4}^{(2)}=$ $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-2(d+1)\vec{q}^{2}J_{31}^{B}+2\vec{q}^{4}J_{32}^{B}+(d^{2}-1)J_{41}^{B}-2(d+1)\vec{q}^{2}J_{42}^{B}+\vec{q}^{4}J_{43}^{B}],$ (28)
$\mathcal{M}_{c1.5}^{(2)}=$ $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A[-\vec{q}^{2}J_{21}^{R}+\vec{q}^{4}J_{22}^{R}-2(d+1)\vec{q}^{2}J_{31}^{R}+2\vec{q}^{4}J_{32}^{R}+(d^{2}-1)J_{41}^{R}-2(d+1)\vec{q}^{2}J_{42}^{R}+\vec{q}^{4}J_{43}^{R}],$ (29)
@@
In the above amplitudes (19)-(29), the loop functions $J^{a/b}_{ij}(m,\omega)$, $J^{g/h}_{ij}(m,\omega_{1},\omega_{2})$, $J^{F}_{ij}(m_{1},m_{2},q)$, $J^{T/S}_{ij}(m_{1},m_{2},\omega,q)$, and $J^{B/R}_{ij}(m_{1},m_{2},\omega_{1},\omega_{2},q)$ are abbreviated as $J^{a/b}_{ij}$, $J^{g/h}_{ij}$, $J^{F}_{ij}$, $J^{T/S}_{ij}$, and $J^{B/R}_{ij}$, respectively. The variables $m,m_{1},m_{2}$ denote the pion mass, while the mass-splitting-dependent variables $\omega_{1}$ and $\omega_{2}$ for every amplitudes are collected in Tables 1-2. Here we define $\delta_{1}=M_{D^{*}}-M_{D}$ and $\delta_{2}=M_{\bar{B}^{*}}-M_{\bar{B}}$ as the $D^{*}$-$D$ and $\bar{B}^{*}$-$\bar{B}$ mass differences, respectively. We list the definitions of the loop functions in Appendix B, and the calculation procedures of these loop functions follow the Refs. Xu:2017tsr ; Xu:2021vsi .
Figure 1: LO contact, NLO contact and NLO TPE diagrams of the process $D\bar{B}\to D\bar{B}$. The thin solid, double thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$ and a pion, respectively.
Table 1: The coefficients appearing in the contact amplitudes (Eqs.( 19)-(24)).

### v3

@@ Abstract
In this work, we systematically study the interactions of the $S$-wave $D^{(*)}\bar{B}^{(*)}$ systems within the framework of chiral effective field theory in heavy hadron formalism. We calculate the $D^{(*)}\bar{B}^{(*)}$ effective potentials up to next-to-leading order, explore the bound state formations, and investigate the $D^{(*)}\bar{B}^{(*)}$ scattering properties such as scattering rate, scattering length, and effective range. Our results show that all $I=1$ $D^{(*)}\bar{B}^{(*)}$ potentials are repulsive, preventing the formation of bound states, while the $I=0$ potentials are generally attractive. Specifically, we get two important observations: first, the shallow bound state is more likely to exist in the $D\bar{B}[I(J^{P})=0(0^{+})]$ system than in the $D\bar{B}^{*}[I(J^{P})=0(1^{+})]$ system; second, $D^{*}\bar{B}^{*}[I(J^{P})=0(0^{+})]$ and $D^{*}\bar{B}^{*}[I(J^{P})=0(1^{+})]$ systems possess relatively large binding energies and positive scattering lengths, which suggests strong bound state formations in these channels. So the attractions in the $D^{*}\bar{B}^{*}[I=0]$ systems are deeper than those in the $D\bar{B}^{(*)}[I=0]$ systems, thus we strongly recommend the future experiment to search for the $D^{*}\bar{B}^{*}[I=0]$ tetraquark systems. We also investigate the coupled-channel effects on the $J=0,1$ systems and conclude that the inclusion of the coupled channels introduces small but visible influences. In addition, we also investigate the dependencies of the $D\bar{B}^{(*)}$ binding energies on the contact low-energy coupling constants.
@@ I Introduction
The study of exotic hadrons has become a vibrant frontier in particle physics, offering profound insights into the nonperturbative regime of strong interactions. In recent years, LHCb Collaboration has made groundbreaking discoveries that have significantly advanced our understanding of multiquark states. In 2020, LHCb reported the observation of two resonances, $X_{0}(2900)$ and $X_{1}(2900)$, in the $D^{-}K^{+}$ invariant mass spectrum of the $B^{+}\to D^{+}D^{-}K^{+}$ decay process LHCb:2020bls ; LHCb:2020pxc . These states, with quantum numbers $J^{P}=0^{+}$ and $J^{P}=1^{+}$, respectively, were identified as the first tetraquark candidates with the exotic quark content $ud\bar{s}\bar{c}$. This discovery was followed in 2021 by the observation of the double-charm tetraquark state $T^{+}_{cc}$ in the $D^{0}D^{0}\pi^{+}$ invariant mass spectrum LHCb:2021vvq ; LHCb:2021auc . Shortly thereafter, LHCb reported two additional open heavy-flavor tetraquark candidates, $T_{c\bar{s}0}^{a}(2900)^{++}$ and $T_{c\bar{s}0}^{a}(2900)^{0}$, in the $D_{s}^{+}\pi^{\pm}$ final states of the $B^{+}\to D^{-}D^{+}_{s}\pi^{+}$ and $B^{0}\to\bar{D}^{0}D^{+}_{s}\pi^{+}$ decays LHCb:2022sfr ; LHCb:2022lzp . These states are believed to have the quark content $c\bar{s}u\bar{d}$.
@@
Doubly heavy dimeson systems have attracted significant theoretical interest due to their potential to form stable or quasistable molecular states. For instance, Li et al. explored the possibility of deuteronlike molecular states in $D^{(*)}D^{(*)}$, $\bar{B}^{(*)}\bar{B}^{(*)}$, and $D^{(*)}\bar{B}^{(*)}$ systems, identifying several promising candidates Li:2012ss . QCD sum rules have been applied to study exotic open-flavor tetraquark states such as $bc\bar{q}\bar{q}$, $bc\bar{s}\bar{s}$, $qc\bar{q}\bar{b}$, and $sc\bar{s}\bar{b}$, revealing that some of these states lie below the $D^{(*)}B^{(*)}$ and $D_{s}^{(*)}B_{s}^{(*)}$ thresholds, making them susceptible to decay via fall-apart mechanisms Chen:2013aba . The color-magnetic interaction model has also been used to calculate mass splittings for $qq\bar{Q}\bar{Q}$ tetraquark states Luo:2017eub , while the molecular picture has been employed to investigate the binding energies of $T_{cc}^{+}$ and its bottom and strange partners Deng:2021gnb . Additionally, the complex scaling method has been utilized to study doubly heavy tetraquark bound and resonant states Wu:2024zbx .
The stability of doubly heavy tetraquarks, particularly those involving bottom and charm quarks, remains a topic of active debate. While some studies suggest that states such as $b\bar{u}\bar{d}, bb$$b\bar{u}\bar{s}$, and $bb\bar{d}\bar{s}$ are stable against strong decay Eichten:2017ffp , others argue that bottom-charm tetraquarks may not be stable Silvestre-Brac:1993zem ; Ebert:2007rn ; Park:2018wjk ; Lu:2020rog ; Braaten:2020nwp ; Kim:2022mpa ; Song:2023izj . The production mechanisms and detection prospects for these states at colliders have also been explored, with estimates suggesting significant production cross sections for double-bottom and mixed bottom-charm tetraquarks at the LHC Ali:2018xfq .
Recent lattice QCD studies have provided further insights into the properties of doubly heavy tetraquarks. Meinel et al. found no evidence for QCD-stable $\bar{b}\bar{c}ud$ tetraquark states Meinel:2022lzo , while Alexandrou et al. and Padmanath et al. reported shallow bound states for isoscalar $\bar{b}\bar{c}ud$ systems Alexandrou:2023cqg ; Padmanath:2023rdu . These findings, along with the identification of subthreshold poles in the $D\bar{B}$ scattering amplitude Radhakrishnan:2024ihu , significantly advanced our understanding of the weak decay behaviors of $D\bar{B}$ molecules Liu:2024lmv .
@@ II.1 Lagrangians at the leading order
where $D_{a}$, $D_{b}$, $E_{a}$, and $E_{b}$ are independent low-energy coupling constants (LECs), and $\tau^{\alpha}$ represents the Pauli matrix in isospin space. The heavy meson doublet $(D,D^{*})$ or $(\bar{B},\bar{B}^{*})$ is described by the $H$ field in the heavy hadron formalism Xu:2017tsr ; Xu:2021vsi ; Wang:2018atz :
@@
where $v=(1,0,0,0)$ is the four-velocity of the heavy mesons. For the $D^{*}\bar{B}^{*}$ system, the scattering amplitudes receive contributions from the one-pion-exchange interactions at LO. The Lagrangian for the LO $D^{*}D^{*}\pi$ vertex or $\bar{B}^{*}\bar{B}^{*}\pi$ vertex is given by Xu:2017tsr ; Xu:2021vsi ; Wang:2018atz
@@
where $\Delta$ in the last term represents the mass splitting between $D$ and $D^{*}$ (or $\bar{B}$ and $\bar{B}^{*}$), $\Gamma_{\mu}$ is the chiral connection, $u_{\mu}$ is the axial vector current, and $\xi=\exp(i\phi/2f)$. Here, $f$ is the bare pion decay constant, and the pion field $\phi$ is
@@ II.2 Lagrangians at next-to-leading order
$\mathcal{L}_{4H}^{(2,h)}=$ $D_{a}^{h}\mathrm{Tr}[H\gamma_{\mu}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\bar{H}]\mathrm{Tr}(\chi_{+})$
$+D_{b}^{h}\mathrm{Tr}[H\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}(\chi_{+})$
$+E_{a}^{h}\mathrm{Tr}[H\gamma_{\mu}\tau^{a}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\tau_{a}\bar{H}]\mathrm{Tr}(\chi_{+})$
$+E_{b}^{h}\mathrm{Tr}[H\gamma_{\mu}\gamma_{5}\tau^{a}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\gamma_{5}\tau_{a}\bar{H}]\mathrm{Tr}(\chi_{+}),$ (6)
$\mathcal{L}_{4H}^{(2,v)}=$ $\{D_{a1}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}(v\cdot D\bar{H})]\mathrm{Tr}[H\gamma^{\mu}\bar{H}]$
$+D_{a2}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}\bar{H}]\mathrm{Tr}[(v\cdot DH)\gamma^{\mu}\bar{H}]$
$+D_{a3}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}(v\cdot D\bar{H})]$
$+D_{a4}^{v}\mathrm{Tr}[((v\cdot D)^{2}H)\gamma_{\mu}\bar{H}]\mathrm{Tr}[H\gamma^{\mu}\bar{H}]$
@@
$+E_{b1}^{v}\mathrm{Tr}[(v\cdot DH)\gamma_{\mu}\gamma_{5}\tau^{a}(v\cdot D\bar{H})]\mathrm{Tr}[H\gamma^{\mu}\gamma_{5}\tau_{a}\bar{H}]$
$+\cdots\}+\mathrm{H.c.},$ (7)
$\mathcal{L}_{4H}^{(2,q)}=$ $\{D_{1}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}(D^{\nu}\bar{H})]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}\bar{H}]$
$+D_{2}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[(D^{\nu}H)\gamma_{\nu}\gamma_{5}\bar{H}]$
$+D_{3}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}(D^{\nu}\bar{H})]$
$+D_{4}^{q}\mathrm{Tr}[(D^{\mu}D^{\nu}H)\gamma_{\mu}\gamma_{5}\bar{H}]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}\bar{H}]$
$+E_{1}^{q}\mathrm{Tr}[(D^{\mu}H)\gamma_{\mu}\gamma_{5}\tau^{a}(D^{\nu}\bar{H})]\mathrm{Tr}[H\gamma_{\nu}\gamma_{5}\tau_{a}\bar{H}]$
$+\cdots\}+\mathrm{H.c.},$ (8)
@@
The low-energy constants appearing in the above equations are split into finite and infinite parts. The infinite parts are used to cancel the divergences of the loop contributions at NLO. The finite parts also contribute, but are not determined here due to the lack of data input. Therefore, we just discard them.
After the scattering amplitudes for the $D^{(*)}\bar{B}^{(*)}$ systems are calculated using the above Lagrangians, the effective potentials in momentum space can be obtained via the relation,
@@
where the factor 1/4-1/4 comes from the Breit approximation. To understand the interactions and further investigate whether the $D^{*}\bar{B}^{(*)}$ systems can form stable molecular states, we need the potentials in coordinate space. Fourier transformations of the momentum-space potentials allow for their derivation,
@@
Simultaneously, according to the effective potentials in momentum space, we can also solve the Lippmann-Schwinger equations to calculate the partial-wave two-body scattering amplitudes $T$,
@@
Considering the $S$-wave, in the low-energy limit the $k\rm{cot}\delta_{0}(k)$ can be expanded in powers of $k^{2}$,
@@
where $a$ is the scattering length and $r_{0}$ is the effective range. The ellipsis represents higher-order terms in $k$. The $S$-wave scattering cross section is proportional to the phase shift,
@@ III.1 $D\bar{B}$ systems
Next, we consider NLO contributions illustrated in Fig. 1. There are two types of NLO diagrams. The diagrams in Fig. 1 represent one-loop corrections to the LO contact interaction, and the corresponding amplitudes are given by,
@@
Using the Lagrangians in Eq. (3), the amplitudes for the TPE diagrams in Fig. 1 can be written as
$\mathcal{M}_{c1.1}^{(2)}=$ $-4\frac{1}{f^{4}}[A_{1}(q_{0}^{2}J_{0}^{F}+J_{22}^{F})$
$-A_{15}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})$
$-A_{51}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})$
$+A_{5}(q_{0}^{2}J_{0}^{F}+2q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})],$ (25)
$\mathcal{M}_{c1.2}^{(2)}=$ $-4i\frac{g_{2}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}$
$+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]$
$-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}-(d-1)q_{0}J_{21}^{T}$
$+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}$
$+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]\},$ (26)
$\mathcal{M}_{c1.3}^{(2)}=$ $-4i\frac{g_{2}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}$
$+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]$
$-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}-(d-1)q_{0}J_{21}^{T}$
$+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}-(d-1)q_{0}J_{31}^{T}$
$+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}-(d-1)J_{34}^{T}]\},$ (27)
$\mathcal{M}_{c1.4}^{(2)}=$ $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-2(d+1)\vec{q}^{2}J_{31}^{B}$
$+2\vec{q}^{4}J_{32}^{B}+(d^{2}-1)J_{41}^{B}$
$-2(d+1)\vec{q}^{2}J_{42}^{B}+\vec{q}^{4}J_{43}^{B}],$ (28)
$\mathcal{M}_{c1.5}^{(2)}=$ $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A[-\vec{q}^{2}J_{21}^{R}+\vec{q}^{4}J_{22}^{R}-2(d+1)\vec{q}^{2}J_{31}^{R}$
$+2\vec{q}^{4}J_{32}^{R}+(d^{2}-1)J_{41}^{R}$
$-2(d+1)\vec{q}^{2}J_{42}^{R}+\vec{q}^{4}J_{43}^{R}],$ (29)
@@
In the above amplitudes (19)-(29), the loop functions $J^{a/b}_{ij}(m,\omega)$, $J^{g/h}_{ij}(m,\omega_{1},\omega_{2})$, $J^{F}_{ij}(m_{1},m_{2},q)$, $J^{T/S}_{ij}(m_{1},m_{2},\omega,q)$, and $J^{B/R}_{ij}(m_{1},m_{2},\omega_{1},\omega_{2},q)$ are abbreviated as $J^{a/b}_{ij}$, $J^{g/h}_{ij}$, $J^{F}_{ij}$, $J^{T/S}_{ij}$, and $J^{B/R}_{ij}$, respectively. The variables $m,m_{1}$, and $m_{2}$ denote the pion mass, while the mass-splitting-dependent variables $\omega_{1}$ and $\omega_{2}$ for every amplitude are collected in Tables 1 and 2. Here, we define $\delta_{1}=M_{D^{*}}-M_{D}$ and $\delta_{2}=M_{\bar{B}^{*}}-M_{\bar{B}}$ as the $^{*}-DD$ and $\bar{B}^{*}$-$\bar{B}$ mass differences, respectively. We list the definitions of the loop functions in Appendix B, and the calculation procedures of these loop functions follow the Refs. Xu:2017tsr ; Xu:2021vsi .
Figure 1: LO contact ($a1$), NLO contact (a1.1a1.6a1.1-a1.6) and NLO TPE (c1.1c1.5c1.1-c1.5) diagrams of the process $D\bar{B}\to D\bar{B}$. The thin solid, double-thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$, and a pion, respectively.
Table 1: The coefficients appearing in the contact amplitudes [Eqs.( 19)-(24)].